# Energetic Resilience of Linear Driftless Systems

### v2

@@ IV-B Energetic Resilience Metrics
Using (16), (33) and (10),
@@
Note that
@@
where $L\coloneqq\lambda_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}\right)$ and we use the Rayleigh inequality on the symmetric matrices $B_{c}^{\dagger T}B_{c}^{\dagger}$ and $B^{\dagger T}B^{\dagger}$. Next,
$\sup_{\|x_{0}\|_{2}\geq R}2t_{f}\frac{\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|}{\|B^{\dagger}x_{0}\|_{2}^{2}}$
$\leq 2t_{f}\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}\sup_{\|x_{0}\|_{2}\geq R}\frac{\|x_{0}\|_{2}}{\|B^{\dagger}x_{0}\|_{2}^{2}}$
$\leq 2t_{f}\frac{\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}}{R\lambda_{\min}\left(B^{\dagger T}B^{\dagger}\right)},$
where the first inequality follows from the Cauchy-Schwarz inequality and the second inequality follows from the Rayleigh inequality on $B^{\dagger T}B^{\dagger}$. Finally,
$\sup_{\|x_{0}\|_{2}\geq R}t_{f}^{2}\frac{\|B_{uc}\|_{2}^{2}+1}{\|B^{\dagger}x_{0}\|_{2}^{2}}\leq t_{f}^{2}\frac{\|B_{uc}\|_{2}^{2}+1}{R^{2}\lambda_{\min}\left(B^{\dagger T}B^{\dagger}\right)},$
since $B^{\dagger T}B^{\dagger}$ is symmetric. Substituting the above three inequalities into (38), we obtain
$r_{M}(t_{f},R)$
$\geq\frac{R^{2}\lambda_{\min}\left(B^{\dagger T}B^{\dagger}\right)}{R^{2}L+2Rt_{f}\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}+t_{f}^{2}\left(\|B_{uc}\|_{2}^{2}+1\right)},$
concluding the proof. ∎
In the following section, we present an example demonstrating the use of these metrics in quantifying the maximal additional energy used by the malfunctioning system compared to the nominal system.
(a) The difference of energies, and the additive metric.
(b) The ratio of energies, and the multiplicative metric.
Figure 1: Illustration of energetic resilience metrics as a function of distance of the initial condition from the origin.
V Simulation Example
We now illustrate the use of the energetic resilience metrics on a driftless model of an underwater robot. We consider a minor modification of the model considered by [13]:
@@ V Simulation Example
where the input matrix is $B$. Here, $x$ and $y$ are the coordinates of the robot. The engine $u_{1}$ acts primarily along the $x$-direction, with a small bias in the $y$-direction. $u_{2}$ and $u_{3}$ are engines at a $45^{\circ}$ angle to the $x$- and $y$-directions, opposing each other. The model in [13] heavily biased the engine $u_{1}$ towards the $x$-direction, and the corresponding entry in the input matrix was $10$. We modify the entry to $2$, reducing the bias in favor of a more equitable distribution of the effect of each input.
@@
We assume that a malfunction causes the system to lose authority over $u_{3}$. Then, $B_{c}=\begin{bmatrix}2&\phantom{-}1\\ 0.2&-1\end{bmatrix}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ B_{uc}=\begin{bmatrix}1\\ 1\end{bmatrix}.$ In what follows, for notational simplicity, we denote $E_{N}\equiv E_{N}^{*}(x_{0},t_{f})$, $\overline{E}_{M}\equiv\overline{E}_{M}(x_{0},t_{f})$ and $E_{M}^{+}\equiv E_{M}^{+}(x_{0},t_{f},u_{uc})$. We now compare the nominal energy $E_{N}$ (16) to the total energy $E_{M}^{+}$ in (30) and worst-case total energy $\overline{E}_{M}$ in (33). In particular, we demonstrate the accuracy of the energetic resilience bounds (35) and (37) in quantifying this comparison.
We fix $t_{f}=10$ and vary the distance $R$ of the initial condition $x_{0}$ from the origin. Figure 1(a) shows the difference between the worst-case total energy $\overline{E}_{M}$ and the nominal energy $E_{N}$ as a function of the distance $R$ of the initial condition $x_{0}$ from the origin. We also plot the upper bound on $r_{A}(t_{f},R)$ from (35) as a function of $R$, and note that this is a useful upper bound characterizing the additional energy used by the malfunctioning system.
@@
Similarly, Fig. 1(b) shows the ratio of the nominal energy $E_{N}$ and the total energy $E_{M}^{+}$ as a function of $R$. Here, we test a large class of uncontrolled inputs, including constant, full-amplitude inputs of the form (32) and different classes of low- and high-frequency sinusoids. We also plot the lower bound on $r_{M}(t_{f},R)$ from (37) as a function of $R$, shown in the thick red line. It is evident that $r_{M}(t_{f},R)$ bounds the ratio of nominal and augmented malfunctioning energies from below. Furthermore, the thick blue line represents the ratio $E_{N}/E_{M}^{+}$ for the worst-case uncontrolled input in (32), where $E_{M}^{+}=\overline{E}_{M}$. The bound on $r_{M}(t_{f},R)$ is a reasonable approximation for this ratio, especially for initial conditions closer to the origin. For instance, when $R=10$, we have $r_{M}(t_{f},R)\approx 0.03$, while $E_{N}/\overline{E}_{M}\approx 0.05$. The bound on $r_{M}(t_{f},R)$ indicates that at most $1/0.03=33.33$ times more energy is used by the actuators to achieve finite-time regulation when a loss of control authority of the form described above occurs. The actual ratio indicates that around $1/0.05\approx 20$ times the energy is required to achieve the task. The metric $r_{M}(t_{f},R)$ is thus a useful quantity to characterize the maximal additional energy required to achieve a task when control authority is lost over a subset of actuators.
@@
If we considered the model used by [13], where the effect of $u_{1}$ on $x$ is weighted more, these metrics would be relatively less useful in characterizing the additional energy required, in comparison to the model in (39). The heavier weight contributes to poor conditioning of the matrices $B^{\dagger T}B^{\dagger}$ and $B_{c}^{\dagger T}B_{c}^{\dagger}$, resulting in more conservative bounds in (35) and (37) than in the example presented here.
VI Conclusions and Future Work
@@
In this paper, our objective was to quantify the maximal additional energy used by a system which loses control authority over a subset of actuators, compared to the nominal system with full control authority. To this end, we considered the special case of linear driftless systems and introduced additive and multiplicative energetic resilience metrics. These metrics compare the nominal and worst-case total energies to achieve finite-time regulation. Deriving the nominal and worst-case total energies for this task used a technical lemma proved using the calculus of variations. We also obtained the optimal final time that minimizes the malfunctioning energy for a given uncontrolled input. When considering the special case of losing control authority over one actuator, we obtained an exact expression for the worst-case total energy, allowing us to obtain bounds on the energetic resilience metrics. A simulation example on a model of an underwater robot demonstrated the applicability of these metrics in characterizing the additional energy used by a malfunctioning system.
Future work involves examining general linear and nonlinear systems. Obtaining energy-optimal control signals in these cases is not a straightforward task due to the input constraints $u\in\operatorname*{\mathcal{U}}$ and $u_{c}\in\operatorname*{\mathcal{U}_{c}}$. Other complexities, such as systems with time delays and systems whose general structure is uncertain due to unmodeled dynamics, are also to be examined.

### v3

@@ 4.2 Energetic Resilience Metric
We provide a brief sketch of the proof to conserve space. Using (15), (27) and (9),
@@
Let $L\coloneqq\lambda_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}\right)$. Using the Rayleigh inequality on $B_{c}^{\dagger T}B_{c}^{\dagger}$ and $B^{\dagger T}B^{\dagger}$ and the Cauchy-Schwarz inequality, each term in the supremum can be bounded as follows:
@@
$\sup_{\|x_{0}\|_{2}\geq R}2t_{f}\frac{\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|}{\|B^{\dagger}x_{0}\|_{2}^{2}}$ $\leq 2t_{f}\frac{\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}}{R\lambda_{\min}\left(B^{\dagger T}B^{\dagger}\right)},$
$\sup_{\|x_{0}\|_{2}\geq R}t_{f}^{2}\frac{\|B_{uc}\|_{2}^{2}+1}{\|B^{\dagger}x_{0}\|_{2}^{2}}$ $\leq t_{f}^{2}\frac{\|B_{uc}\|_{2}^{2}+1}{R^{2}\lambda_{\min}\left(B^{\dagger T}B^{\dagger}\right)}.$
Substituting in (30), the result follows.
In the following section, we present an example demonstrating the use of this metric in quantifying the maximal additional energy used by the malfunctioning system compared to the nominal system.
5 Simulation Example
We now illustrate the use of the energetic resilience metric on a driftless model of an underwater robot. We consider a minor modification of the model considered by Bouvier and Ornik (2020):
@@ 5 Simulation Example
where the input matrix is $B$. Here, $x$ and $y$ are the coordinates of the robot. The engine $u_{1}$ acts primarily along the $x$-direction, with a small bias in the $y$-direction. $u_{2}$ and $u_{3}$ are engines at a $45^{\circ}$ angle to the $x$- and $y$-directions, opposing each other. The model in Bouvier and Ornik (2020) heavily biased the engine $u_{1}$ towards the $x$-direction, and the corresponding entry in the input matrix was $10$. We reduce this bias by modifying the entry to $2$.
@@
We assume that a malfunction causes the system to lose authority over $u_{3}$. Then, $B_{c}=\begin{bmatrix}2&\phantom{-}1\\ 0.2&-1\end{bmatrix}~{}\text{ and }~{}B_{uc}=\begin{bmatrix}1\\ 1\end{bmatrix}.$ In what follows, for notational simplicity, we denote $E_{N}\equiv E_{N}^{*}(x_{0},t_{f})$, $\overline{E}_{M}\equiv\overline{E}_{M}(x_{0},t_{f})$ and $E_{M}^{+}\equiv E_{M}^{+}(x_{0},t_{f},u_{uc})$. We now compare the nominal energy $E_{N}$ (15) to the total energy $E_{M}^{+}$ from (20) and the definition (7), and worst-case total energy $\overline{E}_{M}$ (27). In particular, we demonstrate the accuracy of the energetic resilience bound (29) in quantifying this comparison. We fix $t_{f}=10$ and vary the distance $R$ of the initial condition $x_{0}$ from the origin.
Figure 1: Ratio of nominal and malfunctioning energies compared to the resilience metric. The thinner dashed lines plot $E_{M}^{+}-E_{N}$ for various uncontrolled inputs.
@@
Fig. 1 shows the ratio of the nominal energy $E_{N}$ and the total energy $E_{M}^{+}$ as a function of $R$ in the thin dashed lines. Here, we test a large class of uncontrolled inputs, including constant, full-amplitude inputs of the form (28) and different classes of low- and high-frequency sinusoids. We also plot the lower bound on $r_{M}(t_{f},R)$ from (29) as a function of $R$, shown in the thick red line. It is evident that $r_{M}(t_{f},R)$ bounds the ratio of nominal and augmented malfunctioning energies from below. Furthermore, the thick blue line represents the ratio $E_{N}/E_{M}^{+}$ for the worst-case uncontrolled input in (28), where $E_{M}^{+}=\overline{E}_{M}$. This ratio is calculated using (15) and (27). The bound on $r_{M}(t_{f},R)$ is a reasonable approximation for this ratio, especially for initial conditions closer to the origin. For instance, when $R=10$, we have $r_{M}(t_{f},R)\approx 0.03$, while $E_{N}/\overline{E}_{M}\approx 0.05$. The bound on $r_{M}(t_{f},R)$ indicates that at most $1/0.03=33.33$ times more energy is used by the actuators to achieve finite-time regulation when a loss of control authority of the form described above occurs. The actual ratio indicates that around $1/0.05\approx 20$ times the energy is required to achieve the task. The metric $r_{M}(t_{f},R)$ is thus a useful quantity to characterize the maximal additional energy required to achieve a task when control authority is lost over a subset of actuators.
@@
If we considered the model used by Bouvier and Ornik (2020), this metric would be relatively less useful in characterizing the additional energy required, in comparison to the model in (31). The heavier weight of $u_{1}$ on $x$ contributes to poor conditioning of the matrices $B^{\dagger T}B^{\dagger}$ and $B_{c}^{\dagger T}B_{c}^{\dagger}$, resulting in a more conservative bound in (29) than in the example presented here.
6 Conclusions and Future Work
@@
In this paper, our objective was to quantify the maximal additional energy used by a system which loses control authority over a subset of actuators, compared to a system with no such malfunction. We thus considered the special case of linear driftless systems and introduced an energetic resilience metric comparing the nominal and worst-case total energies to achieve finite-time regulation. Deriving the nominal and worst-case total energies for this task used a technical lemma proved using the calculus of variations. When considering the special case of losing control authority over one actuator, we obtained an exact expression for the worst-case total energy, allowing us to obtain a bound on the resilience metric. A simulation example on a model of an underwater robot demonstrated the applicability of this metric in characterizing the additional energy used by a malfunctioning system. Our ongoing work involves examining resilience of general nonlinear systems.